OpenStudy (anonymous):

Check work please! :D Find the derivative of f(x) = 5x + 9 at x = 2. I got 0

OpenStudy (anonymous):

@zepdrix @nincompoop @ganeshie8

OpenStudy (kainui):

hey there, can you show me how you calculated the derivative? It appears to be wrong. If you think about it intuitively, the derivative is the slope, correct? So what is the slope of a line at any point? You can actually compare your answer to what you would have gotten in algebra! Good old, y=mx+b you know?

OpenStudy (nincompoop):

show your work please

OpenStudy (anonymous):

wrong

OpenStudy (nincompoop):

the trick is to do the derivative first, and then plug the values later

OpenStudy (akashdeepdeb):

f(x) = 5x + 9 f'(x) = \(\frac{d}{dx} (5x) + \frac{d}{dx} 9\) = \(5 \frac{dx}{dx} + \frac{d}{dx} 9\) = 5 * 1 + 0 [Because, the derivative of a constant = 0] = 5 So, f'(x) or the derivative of the function f(x) is not affected by which x the derivative is calculated. It is constant. That is, it has a constant rate of increase for all x which is '5'. Getting this? :)

OpenStudy (nincompoop):

I was hoping she would show her work.. but thanks AKAdudesheep

OpenStudy (anonymous):

my bad :( I thought the second to last step was *0 not +0... Maybe I should fix my handwriting @AkashdeepDeb

OpenStudy (nincompoop):

next time, let the person go through her mistakes so she can fix them more precisely this is a course that thinking is involved :)

OpenStudy (kainui):

Yeah for sure. Calculus is definitely a good and awesome turning point in math for multiple reasons. =P